Torrey painted some signs yellow and the rest red. The number of yellow signs was 5 less than 3 times the number of red signs. Furthermore, 6 times the number of red signs was 70 less than 10 times the number of yellow signs. How many signs were yellow and how many were red?
Answer by Paul(988)

Red signs =x
Yellow signs = 3x-5
Eqaution:
6(x)=10(3x-5)-70
6x=30x-50-70
-24x=-120
x=5
3(5)-5=10
Hence, their were 5 red sings and 10 yellow signs.
Paul.